Suppose you know that a company’s stock currently sells for $56 per share and the required return on the stock is 10 percent. You also know that the total return on the stock is evenly divided between a capital gains yield and a dividend yield. If it’s the company’s policy to always maintain a constant growth rate in its dividends, what is the current dividend per share? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)

  Current dividend per share $   
Answer #1

Capital gains yield and dividend yield is evenly divided

required rate = capital gains yield + dividends yield

10% = 5% + 5%

Therefore, dividend yield is 5%

Dividend yield = D1 / price

0.05 = D1 / 56

D1 = 2.8

Current dividend per share = D1 / (1 + growth)n

Current dividend per share = 2.8 / (1 + 0.05)1

Current dividend per share = $2.67